OpenStudy (anonymous):

(tan²θ/sin²θ) - tan²θ possible answers are: A. 1 B. 0 C. sec²θ D. cot²θ E. cscθ

OpenStudy (paounn):

1.\[{sen^2 \theta \over \cos^2 \theta }* {1 \over sen^2 \theta } -{ sen^2 \theta \over \cos^2 \theta} = { 1 \over \cos^2 \theta} -{ sen^2 \theta \over \cos^2 \theta} = {{1 - sen^2 \theta} \over {\cos^2 \theta} } = {\cos^2 \theta \over \cos^2 \theta} = 1\]